OpenStudy (anonymous):

Owen scored ninety-two points on a one hundred-point test and twenty-seven on a thirty-point quiz. On which did he get a larger fraction of the total points correct? Show your work.

OpenStudy (lucaz):

\[\huge{\frac{92}{100}~and~\frac{27}{30}}\]

OpenStudy (anonymous):

no i have to make it smaller

OpenStudy (anonymous):

ok

OpenStudy (lucaz):

simplifying \[\huge{\frac{23}{25}~and~\frac{9}{10}}\] to compare them you need equal denominators \[\huge{\frac{2}{2}\times\frac{23}{25}~and~\frac{5}{5}\times\frac{9}{10}}\]

OpenStudy (lucaz):

\[\huge{compare}~\frac{46}{50}~and~\frac{45}{50}\]

OpenStudy (anonymous):

thanks freind